OpenStudy (anonymous):

Use a sum or difference formula to find the exact value of the given expression. sin (− π/12)

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Hey :) Ooo offline >.< darn I just missed y a

zepdrix (zepdrix):

Writing the fraction in terms of two other simpler angles can be a little tricky. This is how I would approach it:\[\Large\rm -\frac{\pi}{12}=\frac{3\pi}{12}-\frac{4\pi}{12}\] \[\Large\rm =\frac{\pi}{4}-\frac{\pi}{3}\]

zepdrix (zepdrix):

\[\Large\rm \sin\left(-\frac{\pi}{12}\right)=\sin\left(\frac{\pi}{4}-\frac{\pi}{3}\right)\]And just apply your Difference Formula from there, yes? :)